Mr.P has 4/3 money as Mr.Q, but Mr.Q has 5/4 money as R . Altogether they have Rs.94. How much have each of them.
Answer by ptaylor(2198)

Let x=amt of money that R has
Then Q has (5/4)x
And P has (4/3)(5/4)x
Sooooo
(4/3)(5/4)=20/12=5/3
x+(5/4)x+(5/3)x=94 multiply each term by 12 to get rid of fractions
12x+15x+20x=1128
47x=1128
x=24 ---R's amt
(5/4)*24=30----Q's amt
(5/3)*24=40----P's amt
Hope this helps--ptaylor